# Correlation between particle spectra and elliptic flow

### v1

@@ Abstract
We introduce a new observable to probe the collective nature of the radial expansion of the quark-gluon plasma. This observable, dubbed $v_{02}(p_{T})$, represents the correlation of the spectrum with elliptic flow, in the same way as the recently measured $v_{0}(p_{T})$ represents the correlation of the spectrum with the transverse momentum per particle. The advantage of $v_{02}(p_{T})$ over $v_{0}(p_{T})$ is that it is measured using a three-particle cumulant, as opposed to a pair correlation, which significantly reduces the sensitivity to nonflow effects. We predict non-trivial differences between $v_{02}(p_{T})$ and $v_{0}(p_{T})$ in semi-central Pb+Pb collisions at the Large Hadron Collider (LHC) on the basis of hydrodynamic simulations. A hint of these differences can be seen in the modification of $p_{T}$ spectra observed by ALICE in event-shape-engineered events a decade ago.
@@ I Introduction
We define $v_{02}(p_{T})$ in Sec. II. We show in Sec. III that it is a differential version of Bożek’s correlator between $[p_{T}]$ and $v_{2}$ Bozek:2016yoj , which has been thoroughly studied at RHIC and LHC ATLAS:2019pvn ; ALICE:2021gxt ; ATLAS:2022dov ; STAR:2024wgy ; CMS:2024rvk ; STAR:2024ngl . In Sec. IV, we evaluate $v_{02}(p_{T})$ in hydrodynamic simulations of mid-central Pb+Pb collisions at LHC energy and compare it with $v_{0}(p_{T})$. The qualitative differences between $v_{02}(p_{T})$ and $v_{0}(p_{T})$ are interpreted in Sec. V within a two-component model, where the spectrum is determined by $[p_{T}]$ and $v_{2}$. In Sec. VI, we discuss how the dependence of spectra on event shapes observed by ALICE ALICE:2015lib can be related to $v_{02}(p_{T})$.
@@ II Definitions
Elliptic flow, $v_{2}$, is defined in every event as the modulus of the second Fourier coefficient of the azimuthal distribution of outgoing hadrons Qiu:2011iv , We denote by $n(p_{T})$ the $p_{T}$ spectrum normalized on an event-by-event basis ATLAS:2025ztg , that is,
@@ III Relation with Bożek’s correlator
One sees that the numerators of Eq. (10) and (11) are identical, while the denominators differ. In Eq. (10), we have chosen to normalize the correlation by the uncorrelated part (product of means), while in Eq. (11) it is normalized by the product of standard deviations. Up to this different normalization, $v_{02}$ and $\rho_{2}$ are equivalent observables. Thus $v_{02}(p_{T})$, which is the differential observable associated with $v_{02}$, can be viewed as a differential version of $\rho_{2}$. [^2] We choose to make $\rho_{2}$ differential in the spectrum. It can also be made differential in elliptic flow Samanta:2023rbn .
@@ IV Predictions from hydrodynamic simulations
Figure 1: Hydrodynamic results for charged hadrons (a), pions (b), kaons (c) and protons (d). Solid red lines: $v_{02}(p_{T})$, defined by Eq. (2), from our hydrodynamic simulation, where the band is the statistical error bar due to the finite number of events. Solid black lines: $v_{0}(p_{T})$ defined by Eq. (3), scaled down by a factor $10$. Dashed lines are the result of the two-component model (Eqs. (19) and (20)), where the response functions $\alpha(p_{T})$ and $\beta(p_{T})$ are evaluated by solving hydrodynamics with smooth initial density profiles (see text).
The values of $v_{02}(p_{T})$ returned by our hydrodynamic simulations are displayed in Fig. 1 for charged and identified hadrons. We also evaluate $v_{0}(p_{T})$, which is displayed for the sake of comparison. One sees that the $p_{T}$ dependences differ for the two observables. While $v_{0}(p_{T})$ changes sign once and is negative at small $p_{T}$, $v_{02}(p_{T})$ changes sign twice and is positive at small $p_{T}$. The trend is more pronounced for heavier particles, which have larger $v_{02}(p_{T})$ at small $p_{T}$. Thus we predict an inversion of the standard mass ordering observed for elliptic flow $v_{2}(p_{T})$ Huovinen:2001cy and for $v_{0}(p_{T})$ ALICE:2025iud .
@@ V Two-component model
We now interpret these qualitative differences between $v_{02}(p_{T})$ and $v_{0}(p_{T})$ within a simple model. We assume that the spectrum is solely determined by $[p_{T}]$ and $v_{2}$. Carrying out a fluctuation decomposition $n(p_{T})=\langle n(p_{T})\rangle+\delta n(p_{T})$, $[p_{T}]=\langle p_{T}\rangle+\delta p_{T}$ and $v_{2}^{2}=\langle v_{2}^{2}\rangle+\delta v_{2}^{2}$ and linearizing in the fluctuations, one obtains [^3] The reason why we consider the square of elliptic flow $v_{2}^{2}$ rather than $v_{2}$ ittself is that $v_{2}$ is, strictly speaking, a complex Fourier coefficient Qiu:2011iv . For symmetry reasons, the modification of a scalar, such as the spectrum, can only appear at quadratic order.
@@
where we have used Eq. (11). Thus we have expressed $v_{02}(p_{T})$ and $v_{0}(p_{T})$ as linear combinations of the response functions $\alpha(p_{T})$ and $\beta(p_{T})$, but with different weights. If $v_{2}$ and $[p_{T}]$ are uncorrelated, $v_{02}(p_{T})$ is proportional to $\beta(p_{T})$ while $v_{0}(p_{T})$ is proportional to $\alpha(p_{T})$: This implies in particular that $v_{0}(p_{T})$ is negative at small $p_{T}$ while $v_{02}(p_{T})$ is positive, as observed in Fig. 1. In practice, $v_{2}$ and $[p_{T}]$ are weakly correlated since $\rho_{2}$ is much smaller than unity. This simple two-component model thus naturally explain the qualitative differences between $v_{02}(p_{T})$ and $v_{0}(p_{T})$ seen in our simulations.
@@
We finally inject these response functions into Eqs. (19) and (20), and evaluate the $p_{T}$-integrated quantities in the right-hand side using the event-by-event simulation presented in Sec. IV. The results are displayed as dashed lines in Fig. 1, The two-component model quantitatively explains the differences between $v_{02}(p_{T})$ and $v_{0}(p_{T})$. The weights of $\alpha(p_{T})$ and $\beta(p_{T})$ are of the same order of magnitude, whereas the weight of $\beta(p_{T})$ in $v_{0}(p_{T})$ is practically negligible. Therefore, the different $p_{T}$ dependence of $v_{02}(p_{T})$ and $v_{0}(p_{T})$ is mostly due to the response coefficient $\beta(p_{T})$, as anticipated. The small discrepancies between the two-component model and the full simulation can be attributed to the fact that we evaluate the response functions using smooth initial conditions. The generic effects of fluctuating initial conditions on observables have long been known Andrade:2008xh . They enhance particle production but reduce elliptic flow $v_{2}(p_{T})$ at large $p_{T}$. More generally, they reduce collective effects at large $p_{T}$. The results displayed in Fig. 1 show that this reduction also applies to $v_{0}(p_{T})$ Parida:2024ckk and $v_{02}(p_{T})$.
@@ VI Relation with event-shape-engineering analyses
The first application of the ESE technique by the ALICE Collaboration, a decade ago, was to hadron spectra ALICE:2015lib . They measured their relative change from small-$q_{2}$ to large-$q_{2}$ Pb+Pb collision events. This amounts to studying the correlation of spectra with elliptic flow, as measured by $v_{02}(p_{T})$ defined in Sec. II. There is no direct quantitative relation between ESE results and $v_{02}(p_{T})$ for two reasons: First, ESE is essentially a qualitative technique. One observes in which direction a given observable changes as $v_{2}$ increases, but one does not extract any quantitative measure of the correlation of that particular observable with $v_{2}$. A second difference is that we define $v_{02}(p_{T})$ using the fraction of particles $n(p_{T})$ in the $p_{T}$ bin, as opposed to the number of particles $N(p_{T})$. This amounts to normalizing the spectra on an event-by-event basis, and such a normalization is not implemented in the ESE analysis. If one replaces $n(p_{T})$ with $N(p_{T})$ in the definition of $v_{02}(p_{T})$, it is shifted by a constant proportional to the correlation between the total multiplicity and $v_{2}^{2}$. With these two differences taken into account, one expects:
@@
where $a$ and $b$ are dimensionless constants, which are independent of $p_{T}$ and particle species, but depend on the bounds on $q_{2}$ which define large-$q_{2}$ and small-$q_{2}$ event samples.
Fig. 2 displays our fit to ALICE data using Eq. (21), where $v_{02}(p_{T})$ is taken from the hydrodynamic calculation in Fig. 1. We obtain a reasonable fit with $a=1.5$ and $b=0.007$. This comparison between theory and data shows that the hydrodynamic calculation simultaneously captures the $p_{T}$ dependences and the relative ordering of pion, kaon, proton results. The parameter $a$ is of order unity, which means that hydrodynamics naturally captures the order of magnitude of the observed effect.
@@ VII Conclusions
We have introduced a new observable $v_{02}(p_{T})$, which probes collective flow through the correlation between hadron spectra and elliptic flow in heavy-ion collisions. It is a non-trivial probe of collectivity, in the sense that it would vanish in the absence of final-state interactions, in the same way as elliptic flow itself. It is a three-particle cumulant and is therefore much less sensitive to nonflow effects than observables based on pair correlations, such as $v_{0}(p_{T})$. This will be useful in particular at high $p_{T}$, where there are large residual nonflow effects from jets, even after applying a rapidity gap.
We predict that the $p_{T}$ dependence of $v_{02}(p_{T})$ differs significantly from that of $v_{0}(p_{T})$ in mid-central Pb+Pb collisions. In particular, while $v_{0}(p_{T})$ for identified hadrons satisfies the usual mass ordering, the order is inverted for $v_{02}(p_{T})$ at small $p_{T}$. This inversion results from a modification of the spectra which is induced specifically by a change of elliptic flow at fixed $[p_{T}]$. One typically expects that in central collisions, where elliptic flow is smaller, the usual mass ordering will be restored. Detailed hydrodynamic calculations are needed in order to predict at which centrality this effect appears, and whether is it also expected for smaller colliding nuclei.
This new observable thus reveals new phenomena, which probe the pressure gradients in the fluid in a more detailed way than existing observables. It can be easily measured in nucleus-nucleus collisions at the LHC and at RHIC, perhaps also in proton-nucleus collisions. In the same way as the $\rho_{2}$ correlator has been proven a sensitive probe of nuclear deformation Giacalone:2020awm ; Jia:2021wbq ; Bally:2021qys ; Jia:2021qyu ; Bally:2023dxi ; Zhao:2024lpc ; Fortier:2024yxs ; Mantysaari:2024uwn , $\alpha$ clusters in light nuclei Zhang:2024vkh and nucleon width Giacalone:2021clp , we expect that its differential version $v_{02}(p_{T})$ will bring new, tight constraints on nuclear structure. Its generalization to other Fourier harmonics, e.g. the correlation between spectra and triangular flow $v_{03}(p_{T})$ Alver:2010gr ) is straightforward. We anticipate a rich upcoming activity on this topic.
@@ Appendix A Experimental analysis
$\langle XYZ\rangle_{c}=\langle XYZ\rangle-\langle XY\rangle\langle Z\rangle-\langle YZ\rangle\langle X\rangle-\langle ZX\rangle\langle Y\rangle+2\langle X\rangle\langle Y\rangle\langle Z\rangle.$ (25)
@@
When analyzing $v_{02}(p_{T})$ for identified hadrons, the only change is that the numerator of Eq. (26) is the numbrer of these identified hadrons. All other quantities ($X$, $Y$ and $N_{C}$ in the denominator of Eq. (26)) should still be evaluated for unidentified charged hadrons. Note that if the detector has uniform acceptance and efficiency, azimuthal symmetry implies $\langle X\rangle=\langle Y\rangle=\langle XZ\rangle=\langle YZ\rangle=0$, so that Eq. (25) simplifies to $\langle XYZ\rangle_{c}=\langle XYZ\rangle-\langle XY\rangle\langle Z\rangle$. If the acceptance is not uniform, one should keep all the terms in Eq. (25).

### v2

@@ Abstract
We introduce a new observable to probe the collective nature of the radial expansion of the quark-gluon plasma. This observable, dubbed $v_{02}(p_{T})$, represents the correlation of the spectrum with elliptic flow, in the same way as the recently measured $v_{0}(p_{T})$ represents the correlation of the spectrum with the transverse momentum per particle. The advantage of $v_{02}(p_{T})$ over $v_{0}(p_{T})$ is that it is measured using a three-particle cumulant, as opposed to a pair correlation, which significantly reduces the sensitivity to nonflow effects. We predict non-trivial differences between $v_{02}(p_{T})$ and $v_{0}(p_{T})$ in semi-central Pb+Pb collisions at the Large Hadron Collider (LHC) on the basis of hydrodynamic simulations. A hint of these differences can be seen in the modification of $p_{T}$ spectra observed by ALICE in event-shape-engineered events.
@@ I Introduction
We define $v_{02}(p_{T})$ in Sec. II. We show in Sec. III that it is a differential version of Bożek’s correlator between $[p_{T}]$ and $v_{2}$ Bozek:2016yoj , which has been thoroughly studied at RHIC and LHC ATLAS:2019pvn ; ALICE:2021gxt ; ATLAS:2022dov ; STAR:2024wgy ; CMS:2024rvk ; STAR:2024ngl . In Sec. IV, we evaluate $v_{02}(p_{T})$ in hydrodynamic simulations of mid-central Pb+Pb collisions at LHC energy and compare it with $v_{0}(p_{T})$. The qualitative differences between $v_{02}(p_{T})$ and $v_{0}(p_{T})$ are interpreted in Sec. V within a two-component model, where the spectrum is determined by $[p_{T}]$ and $v_{2}$. In Sec. VI, we discuss how the dependence of spectra on event shapes observed by ALICE a decade ago ALICE:2015lib can be related to $v_{02}(p_{T})$, and we compare experimental data with hydrodynamic simulations.
@@ II Definitions
Elliptic flow, $v_{2}$, is defined in every event as the modulus of the second Fourier coefficient of the azimuthal distribution of outgoing hadrons Qiu:2011iv . We denote by $n(p_{T})$ the $p_{T}$ spectrum normalized on an event-by-event basis ATLAS:2025ztg , that is,
@@ III Relation with Bożek’s correlator
The numerators of Eq. (10) and (11) are identical, while the denominators differ. In Eq. (10), we have chosen to normalize the correlation by the uncorrelated part (product of means), while in Eq. (11) it is normalized by the product of standard deviations. Up to this different normalization, $v_{02}$ and $\rho_{2}$ are equivalent observables. Thus $v_{02}(p_{T})$, which is the differential observable associated with $v_{02}$, can be viewed as a differential version of $\rho_{2}$. [^2] We choose to make $\rho_{2}$ differential in the spectrum. It can also be made differential in elliptic flow Samanta:2023rbn .
@@ IV Predictions from hydrodynamic simulations
Figure 1: Hydrodynamic results for unidentified charged hadrons (a), pions (b), kaons (c) and protons (d). Solid red lines: $v_{02}(p_{T})$, defined by Eq. (2), from our hydrodynamic simulation, where the band is the statistical error bar due to the finite number of events. Solid black lines: $v_{0}(p_{T})$ defined by Eq. (3), scaled down by a factor $10$. Dashed lines are the result of the two-component model (Eqs. (19) and (20)), where the response functions $\alpha(p_{T})$ and $\beta(p_{T})$ are evaluated by solving hydrodynamics with smooth initial density profiles (see text).
The values of $v_{02}(p_{T})$ returned by our hydrodynamic simulations are displayed in Fig. 1 for unidentified and identified charged hadrons. In our simulations, “unidentified charged hadrons” stands for pions, kaons or protons. We also evaluate $v_{0}(p_{T})$, which is displayed for the sake of comparison. One sees that the $p_{T}$ dependence differs for the two observables. While $v_{0}(p_{T})$ changes sign once and is negative at small $p_{T}$, $v_{02}(p_{T})$ changes sign twice and is positive at small $p_{T}$. The trend is more pronounced for heavier particles, which have larger $v_{02}(p_{T})$ at small $p_{T}$. Thus we predict an inversion of the standard mass ordering observed for elliptic flow $v_{2}(p_{T})$ Huovinen:2001cy and for $v_{0}(p_{T})$ ALICE:2025iud .
@@ V Two-component model
We now interpret these qualitative differences between $v_{02}(p_{T})$ and $v_{0}(p_{T})$ within a simple model. We assume that the spectrum is solely determined by $[p_{T}]$ and $v_{2}$. Carrying out a fluctuation decomposition $n(p_{T})=\langle n(p_{T})\rangle+\delta n(p_{T})$, $[p_{T}]=\langle p_{T}\rangle+\delta p_{T}$ and $v_{2}^{2}=\langle v_{2}^{2}\rangle+\delta v_{2}^{2}$ and linearizing in the fluctuations, one obtains [^3] The reason why we consider the square of elliptic flow $v_{2}^{2}$ rather than $v_{2}$ itself is that $v_{2}$ is, strictly speaking, a complex Fourier coefficient Qiu:2011iv . For symmetry reasons, the modification of a scalar, such as the spectrum, can only appear at quadratic order.
@@
where we have used Eq. (11). Thus we have expressed $v_{02}(p_{T})$ and $v_{0}(p_{T})$ as linear combinations of the response functions $\alpha(p_{T})$ and $\beta(p_{T})$, but with different weights. If $v_{2}$ and $[p_{T}]$ are uncorrelated, $v_{02}(p_{T})$ is proportional to $\beta(p_{T})$ while $v_{0}(p_{T})$ is proportional to $\alpha(p_{T})$: This implies in particular that $v_{0}(p_{T})$ is negative at small $p_{T}$ while $v_{02}(p_{T})$ is positive, as observed in Fig. 1. In practice, $v_{2}$ and $[p_{T}]$ are weakly correlated since $\rho_{2}$ is much smaller than unity. This simple two-component model thus naturally explains the qualitative differences between $v_{02}(p_{T})$ and $v_{0}(p_{T})$ seen in our simulations.
@@
We finally inject these response functions into Eqs. (19) and (20), and evaluate the $p_{T}$-integrated quantities in the right-hand side using the event-by-event simulation presented in Sec. IV. The results are displayed as dashed lines in Fig. 1. The two-component model quantitatively explains the differences between $v_{02}(p_{T})$ and $v_{0}(p_{T})$. The weights of $\alpha(p_{T})$ and $\beta(p_{T})$ are of the same order of magnitude in $v_{02}(p_{T})$, whereas the weight of $\beta(p_{T})$ in $v_{0}(p_{T})$ is practically negligible. Therefore, the different $p_{T}$ dependence of $v_{02}(p_{T})$ and $v_{0}(p_{T})$ is mostly due to the response coefficient $\beta(p_{T})$, as anticipated. The discrepancies between the two-component model and the full simulation can be attributed to the fact that we evaluate the response functions using smooth initial conditions. The generic effects of fluctuating initial conditions on observables have long been known Andrade:2008xh . They enhance particle production but reduce elliptic flow $v_{2}(p_{T})$ at large $p_{T}$. More generally, they reduce collective effects at large $p_{T}$. The results displayed in Fig. 1 show that this reduction also applies to $v_{0}(p_{T})$ Parida:2024ckk and $v_{02}(p_{T})$.
@@ VI Relation with event-shape-engineering analyses
The first application of the ESE technique by the ALICE Collaboration was to hadron spectra ALICE:2015lib . They measured their relative change from small-$q_{2}$ to large-$q_{2}$ Pb+Pb collision events. This amounts to studying the correlation of spectra with elliptic flow, so that this ten-year-old analysis pioneered the study of $v_{02}(p_{T})$. However, the ESE results cannot be directly compared with $v_{02}(p_{T})$ as defined in Sec. II for two reasons: First, the dependence of a given observable on $q_{2}$ is proportional to the correlation of that observable with $v_{2}$, but the proportionality constant depends on the detector. Second, we define $v_{02}(p_{T})$ using the fraction of particles $n(p_{T})$ in the $p_{T}$ bin, as opposed to the number of particles $N(p_{T})$. This amounts to normalizing the spectra on an event-by-event basis, and such a normalization is not implemented in the ESE analysis. If one replaces $n(p_{T})$ with $N(p_{T})$ in the definition of $v_{02}(p_{T})$, it is shifted by an additive constant proportional to the correlation between the total multiplicity and $v_{2}^{2}$. With these two differences taken into account, one expects:
@@
where $a$ and $b$ are dimensionless constants, which are independent of $p_{T}$ and particle species, but depend on analysis details.
Fig. 2 displays our fit to ALICE data using Eq. (21), where $v_{02}(p_{T})$ is taken from the hydrodynamic calculation in Fig. 1. We obtain a reasonable fit with $a=1.5$ and $b=0.007$. This comparison between theory and data shows that the hydrodynamic calculation simultaneously captures the $p_{T}$ dependences and the relative ordering of pion, kaon, proton results. In particular, data seem to support the inverted mass ordering predicted by hydrodynamics at low $p_{T}$. The parameter $a$ is of order unity, which means that hydrodynamics naturally captures the order of magnitude of the observed effect.
@@ VII Conclusions
We have introduced a new observable $v_{02}(p_{T})$, which probes collective flow through the correlation between hadron spectra and elliptic flow in heavy-ion collisions. It is straightforward to measure, and the results can be directly compared to model calculations. In this respect, it represents a significant improvement over the use of ESE events pioneered by ALICE ALICE:2015lib . $v_{02}(p_{T})$ is a non-trivial probe of collectivity, in the sense that it would vanish in the absence of final-state interactions, in the same way as elliptic flow itself. It is a three-particle cumulant and is therefore much less sensitive to nonflow effects than observables based on pair correlations, such as $v_{0}(p_{T})$. This will be useful in particular at high $p_{T}$, where there are large residual nonflow effects from jets, even after applying a rapidity gap.
We predict that the $p_{T}$ dependence of $v_{02}(p_{T})$ differs significantly from that of $v_{0}(p_{T})$ in mid-central Pb+Pb collisions. In particular, while $v_{0}(p_{T})$ for identified hadrons satisfies the usual mass ordering, the order is inverted for $v_{02}(p_{T})$ at small $p_{T}$. This inversion results from a modification of the spectra which is induced specifically by a change of elliptic flow at fixed $[p_{T}]$. One typically expects that in central collisions, where elliptic flow is smaller, the usual mass ordering will be restored. Detailed hydrodynamic calculations are needed in order to predict at which centrality this effect appears, and whether it is also expected for smaller colliding nuclei.
We thus expect that the $p_{T}$ dependence of $v_{02}(p_{T})$ will change qualitatively as a function of centrality and system size. This is a new feature, which is not observed with usual differential observables: Particle spectra Muncinelli:2024izj , $v_{0}(p_{T})$ ATLAS:2025ztg and $v_{n}(p_{T})$ ATLAS:2018ezv are almost centrality independent, once properly rescaled. In hydrodynamics, they depend weakly on centrality and on transport coefficients after rescaling Muncinelli:2024izj ; Parida:2024ckk . $v_{02}(p_{T})$ is different because it results from the interplay between two different phenomena. We have modeled it as a linear combination of two different response functions, whose relative weight depends on centrality. Detailed hydrodynamic calculations are also needed to assess the dependence of each component on transport coefficients and initial conditions.
$v_{02}(p_{T})$ thus reveals new phenomena, which probe the pressure gradients in the fluid in a more detailed way than existing observables. It can be measured at the LHC and at RHIC for various colliding systems. In the same way as the $\rho_{2}$ correlator has been proven a sensitive probe of nuclear deformation Giacalone:2020awm ; Jia:2021wbq ; Bally:2021qys ; Jia:2021qyu ; Bally:2023dxi ; Zhao:2024lpc ; Fortier:2024yxs ; Mantysaari:2024uwn , $\alpha$ clusters in light nuclei Zhang:2024vkh and nucleon width Giacalone:2021clp , we expect that its differential version $v_{02}(p_{T})$ will bring new, tight constraints on nuclear structure. Its generalization to other Fourier harmonics, e.g. the correlation between spectra and triangular flow $v_{03}(p_{T})$ Alver:2010gr , is straightforward. We anticipate a rich upcoming activity on this topic.
@@ Appendix A Experimental analysis
$\langle XYZ\rangle_{c}=\langle XYZ\rangle-\langle XY\rangle\langle Z\rangle-\langle YZ\rangle\langle X\rangle-\langle ZX\rangle\langle Y\rangle+2\langle X\rangle\langle Y\rangle\langle Z\rangle,$ (25)
@@
When analyzing $v_{02}(p_{T})$ for identified hadrons, the only change is that the numerator of Eq. (26) is the number of these identified hadrons. All other quantities ($X$, $Y$ and $N_{C}$ in the denominator of Eq. (26)) should still be evaluated for unidentified charged hadrons. Note that if the detector has uniform acceptance and efficiency, azimuthal symmetry implies $\langle X\rangle=\langle Y\rangle=\langle XZ\rangle=\langle YZ\rangle=0$, so that Eq. (25) simplifies to $\langle XYZ\rangle_{c}=\langle XYZ\rangle-\langle XY\rangle\langle Z\rangle$. If the acceptance is not uniform, one should keep all the terms in Eq. (25).